Previous | Calendar |
June |
2015 |
Next | ||||||||||||||||||||||||||||||
---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|
Sunday | Monday | Tuesday | Wednesday | Thursday | Friday | Saturday | ||||||||||||||||||||||||||||
31JD151 |
1JD1521-4 pm, NWT LTER meeting (INSTAAR) |
2JD153Site Visit (EG, MR, DA): Details EG removed one of the compression interceptometers still lots of snow at the site |
3JD154 |
4JD155SpB Off (half-day) |
5JD156at approx 6:00 MST, LI-6262 flow rate drops (inlet or line partially plugged?) at 11:40 MST, LI-7200 locks up (but analog data ok?) T-storms |
6JD157
|
7JD158
|
8JD159Site Visit (SpB): Details , Photos (30) Span cylinder at 1150 psi. N2 at 1750 psi rebooted LI-7200 (consider changing power supply?) swapped SD cards from cloudCAMeast at 12:15 MST, re-leveled subcanopy cnr2 at 13:30 MST, fixed LI-6262 flow constriction problem (see notes) from 14:07-14:10, swiped krypton hygrometer added section of plywood to platform (to shade the DSM) turned on fan in LI-6262 box CU Malibu Sunny with a few clouds (still lots of snow)
|
9JD160
|
10JD161
|
11JD162
|
12JD163
|
13JD164
|
14JD165
|
15JD166at 15:00 MST, LI-7550/LI-7200 locks up (again!)
|
16JD167
|
17JD168
|
18JD169Site Visit (SpB): Details , Photos (33) Span cylinder at 1110 psi. N2 at 1600 psi Brought 4 new N2 cylinders from MRS to C1 and returned 4 empties (see details) from 11:40-11:59 MST, added 2-micron filter back to the LI-6262 box from 12:20-12:30 MST, rebooted LI-7200 + swapped LI-7550 power supply w/ LI-7500A from 12:55-13:05 MST, Changed profiler (li-6251) dessicant turned on fan in LI-6251 box swapped SD cards from cloudCAMwest, snowCAM CU Suburban Sunny with a few clouds (snow is GONE)
|
19JD170
|
20JD171
|
21JD172Summer Solstice
|
22JD173
|
23JD174
|
24JD175Site Visit (SpB): Details , Photos (23) Span cylinder at 1110 psi. N2 at 1600 psi at 11:02 MST, started ops77: added co-located Ttcs with CSATs (started mark_03.csi) at 9:30 MST, rebooted 2.5m LI-7500A (OP1) swapped SD cards from cloudCAMeast, towerbaseCAM MRS Shuttle (from INSTAAR) Overcast most of day (T-storm in afternoon)
|
25JD176Site Visit (MR): Details change inlet filters on TGA
|
26JD177SpB Off (Sacramento)
|
27JD178
|
28JD179
|
29JD180
|
30JD181
|
1JD182
|
2JD183
|
3JD184
|
4JD185 |